Zeus:

a small company has 10000 shares. Steven owns 50 of these shares. What is Stevens ownership percentage after the split

justus:

50 out of 10000 is 0.5%

Zeus:

Thanks for the help

justus:

No problem!

justus:

Or we can calculate it like this 50/10000 x 100 = 0.5

justus:

@dude Subject: math

jimthompson5910:

Justus has the correct answer. Though the fraction written is before the split happens. If we had a 2-for-1 split, then we have twice as many shares Steven would then own 50*2 = 100 shares out of 2*10000 = 20000 total So, 100/20000 = 0.005 = 0.5% is the same as before. For an n-for-1 split, we can write \(\Large \frac{n*50}{n*10000 } = \frac{50}{10000} = 0.005 = 0.5\%\) where n is some positive whole number (such as n = 2). The equation above indicates it doesn't matter how the split happens. Steven will still retain the same ownership percentage. This is a good thing because if it was the case where stock splits dilute a person's ownership stake, then the company could effectively stock split their way to kicking someone out, in favor of someone else. This would be an unfair practice.

justus:

Great explanation!

jimthompson5910:

Thanks I appreciate the compliments.